the total energy of the satelite revolving the earth is negetive. explain the statement. 
 PLS!!!!!!

Grade:9

2 Answers

vikas askiitian expert
13 years ago

net centripital force = mv2/r               (r is radius of orbit)

centripital force is provided by gravitational force so

 GMm/r2 = mv2/r

  mv2 = GMm/r

 mv2/2 = GMm/2r

   KE = GMm/2r     ................1

potential energy of particle = -GMm/r     ............2

total energy = PE + KE = GMm/r - GMm/2r

                             TE =-GMm/2r

                              TE  < 0

hence proved

sai raghava rajeev penagamuri
13 years ago

IT IS SIMILAR TO THE MOTION OF AN ELECTRON INA ORBIT.WEKNOW THAT THE POTENTIAL ENERGY OF AN ELECTRON IS NEGATIVE AS THE FORCE IS ATTRACTIVE.AND ALSO THE POTENTIAL ENERGY IS DOUBLE OF K.E,TOTAL ENERGY. SO,TOTAL ENERGY IS NEGATIVE IN MAGNITUDE